# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v5

@@ 2.1.3 Code-to-Text
2.1.4 Code-to-Pattern
Code-to-pattern tasks conduct classification on code. Related works are listed in Figure 7.
- Type prediction aims to predict the type of dynamic programming languages such as Python and JavaScript. It has been used as a pretraining objective for code language models (Wang et al., 2022d), where it is often simplified as a binary tagging task to predict which tokens in the code are identifiers (Wang et al., 2021c; e).
- Malware detection is a similar task to defect detection. However, malware differs from other types of vulnerable code in that the bugs therein are malicious - i.e. they are intentionally injected.
- Code reasoning is a recently introduced task for evaluating LLMs, and often comes as a subset of general evaluation benchmarks such as MMLU (Hendrycks et al., 2021b). This task requires the model to reason about the code or algorithms, and answer related questions which are written in multiple-choice format or free-form QA format and may range from conceptual understanding to numerical calculation and complexity analysis.
- Machine code detection is another recently introduced task and aims to predict whether the input code is written by human or generated by machines. Detection of AI generated content (AIGC) has undergone thorough studies in general or specialized domains such as education (Mitchell et al., 2023; Liu et al., 2023k), but it is yet to be examined at scale on code.
2.1.5 Text-to-Text
Text-to-text tasks take text as input, and output text. Related works are listed in Figure 7.
- Document translation is the automatic translation of code-related documents. Since models, datasets, and prompting strategies for machine translation are abundant in NLP (Vaswani et al., 2017; Goyal et al., 2022; He et al., 2023b), we do not go into detail about this task.
@@ 2.1.7 Code LLMs for Low-Resource, Low-Level, and Domain-Specific Languages
This phenomenon also exists in code modeling, as most works listed in Figure 4-7 target the most popular programming languages such as C, Java, and Python. However, a major difference between NLP and SE is that in NLP, low-resource languages are often spoken by few people or even endangered, while in SE low-resource languages are often designed for specific domains and purposes, and thus have an active user community. Verilog, a hardware description language, is one such example, for which code modeling research is quite abundant (Thakur et al., 2023a; b; Lu et al., 2023b; Liu et al., 2023g; Tsai et al., 2023; Thorat et al., 2023; Liu et al., 2023h; Nadimi & Zheng, 2024).
@@ 2.2 Evaluation Metrics
As code generation models became more capable over the years, however, these metrics based on content-overlap have been found to be inadequate (Rozière et al., 2020; Hendrycks et al., 2021a; Austin et al., 2021), since functionally equivalent snippets of code can differ dramatically in their lexical forms. Consequently, researchers have turned their attention to functional correctness. One popular example of such metrics is pass@, proposed by Kulal et al. (2019) and refined by Chen et al. (2021b), which is an unbiased estimator of the model’s chance in passing all unit tests of a program with any of generated samples. This metric can be generalized to pass (Li et al., 2022g), which limits the number of model submissions to but allows filtering by unit tests given in the input from samples.
@@ 2.3 Program Synthesis
Many of the early methods for program synthesis are example-based (Menon et al., 2013), which means they induce programs from input-output examples, often in domain-specific languages (DSLs) such as FlashFill (Devlin et al., 2017a) and Karel222FlashFill is used in Microsoft Excel for string transformation. Karel is a simple programming language for educational purpose. (Bunel et al., 2018), as these languages are usually simple in syntax and structure.
As code generation models became more capable over the years, researchers started to pay attention to program synthesis in general-purpose programming languages as well. Hearthstone (Ling et al., 2016) and CONCODE (Iyer et al., 2018) are two of the early datasets, representing Python and Java respectively. Each example in Hearthstone is the description of a card in the game and its corresponding class implementation, while examples in CONCODE are simply Java methods paired with their natural-language documentation crawled from public GitHub repositories. Synthesizing programs from their corresponding natural language descriptions has since then become a standard practice in program synthesis, and has led to some of the most widely used benchmarks, such as HumanEval (Chen et al., 2021b), which has even been translated into multiple languages (Cassano et al., 2023b; Zheng et al., 2023a; Muennighoff et al., 2024). Some recent benchmarks use general-purpose languages but focus on specific domains, such as data science (Bavishi et al., 2019; Lai et al., 2023) or Jupyter notebooks (Agashe et al., 2019), while several math reasoning benchmarks have also been converted to programming tasks, including MathQA-Python (Amini et al., 2019; Austin et al., 2021) and GSM8K-Python (Cobbe et al., 2021; Chowdhery et al., 2023; Wang et al., 2023f).
@@ 2.4 Repository-Level Coding
Most tasks discussed in Section 2.1 are limited to a single file or even a single function, as cross-file code modeling poses challenges that are beyond the capability of most existing language models. Recently, however, position interpolation techniques (Chen et al., 2023b; Rozière et al., 2023; Peng et al., 2023a) have extended the context window of LLMs to hundreds of thousands of tokens, making it possible to contextualize coding activities within entire repositories. Several works have studied code completion (Shrivastava et al., 2023b; Ding et al., 2022b; Zhang et al., 2023b; Shrivastava et al., 2023a; Phan et al., 2024; Wu et al., 2024) and generation (Liao et al., 2023; Zan et al., 2024) leveraging repository-level context, and corresponding benchmarks have been proposed Liu et al. (2023j); Ding et al. (2023); Li et al. (2024a). Bairi et al. (2023) investigate the more challenging tasks of repository-level API migration and temporal editing, while Jimenez et al. (2023) introduce a related benchmark, SWE-bench.
@@ 3.1 Causal Language Modeling
Unidirectional language models (also known as causal language models333The training objective of such language models is Causal Language Modeling (CLM), but also referred to as Next Token Prediction.) factor the probability of a sentence into the product of each token’s conditional probability with the chain rule. A piece of input text consisting of tokens is modeled as
@@ 3.3 Denoising Objectives
Tay et al. (2023b) argue that span corruption is also closely related to CLM, since one can mask out the whole input text as a single span and train the decoder to generate it autoregressively. Inspired by this relation, they propose UL2, which is the combination of many span corruption objectives that differ in corruption rate and span length. Applying it to both encoder-decoder models and decoder-only models, they find that encoder-decoder models perform better under the same computation budget constraint. Other researches have also found that such encoder-decoder models generally perform better than causal decoder-only models (Wang et al., 2022b; Soltan et al., 2022).
@@ 3.5 Implementation Design
As self-attention alone cannot capture the position information of each input token, the vanilla Transformer adds a non-learnable vector that’s dependent on the absolute position in the input sequence to the embedding of each token, which is known as cosine position embedding. Later works such as GPT and BERT use learnable embeddings instead, while T5 adopts relative position embedding (Shaw et al., 2018), where the information of relative position between query and key is injected in the computation of self-attention instead. A more recent scheme RoPE (Su et al., 2024) multiplies the keys and queries by a position-dependent rotation matrix, and is later shown to enable position interpolation for processing of longer sequences (Chen et al., 2023b; Rozière et al., 2023; Peng et al., 2023a). Alternatively, Press et al. (2022) propose ALiBi, which directly attenuates the attention scores according to the relative position between key and query. RoPE is popularized by its application in PaLM (Chowdhery et al., 2023) and GPT-NeoX (Black et al., 2022), while ALiBi is adopted by BLOOM (Scao et al., 2022). We refer to Zhao et al. (2023) for a survey on position embedding and interpolation in Transformers.
@@ 4 General Language Models for Code
Since language models scaled to hundreds of billions of parameters (Brown et al., 2020; Chowdhery et al., 2023), many of them have demonstrated non-trivial coding capability, even if they are not specifically designed or trained for code. Pioneered by Codex, researchers have also found continual pretraining on code to significantly benefit language models’ performance on code555While some works refer to this process as “finetuning on code”, it is still self-supervised in nature. Thus we choose to adopt the term “extra/additional/continual pretraining” in this work to avoid confusion with supervised in-task finetuning or instruction finetuning..
@@ 4.1 Off-the-Shelf Language Models
Another rising trend is mixture-of-expert models (Lepikhin et al., 2021; Fedus et al., 2022; Du et al., 2022). Notably, Jiang et al. (2024a) recently introduce Mixtral 8x7B, where only 13B parameters are activated for each token during inference, but achieving 40.2 on HumanEval and surpassing much larger dense models such as LLaMA 2. Similarly, Dai et al. (2024) present DeepSeekMoE, where only 2.8B parameters are activated in a 16B model, scoring 26.8 on HumanEval.
@@ 5 Specialized Language Models for Code
As pretrained Transformers such as GPT and BERT achieved remarkable success in natural language processing, such model architectures, learning paradigms, and training objectives were soon adopted by the software engineering community to produce specialized models for code understanding and generation. In this section, we first review common datasets used for pretraining code language models (Section 5.1), and then dive into the complex family of code LMs by their model architecture: encoder-only models (Section 5.2), encoder-decoder models (Section 5.3), decoder-only models (Section 5.4), UniLM (Section 5.5), and diffusion models (Section 5.6). Lastly, in Section 5.7 we also illustrate the current trend of applying more recent techniques in NLP, such as instruction tuning (Wei et al., 2022b; Sanh et al., 2022; Chung et al., 2022) and reinforcement learning (Ouyang et al., 2022) to code processing. An overview of these pretrained models are provided in Table 3.
5.1 Training Dataset for Code
@@
Compared with natural language, another byproduct of scraping code from GitHub is commit histories, which consist of code before commit, code after commit, and a short message describing the commit, which can loosely serve as an instruction for language models. Muennighoff et al. (2024) utilize this feature and construct a 2GB dataset CommitPackFT containing 742K samples of instruction data for code, obviating the need of extensive human labor that’s required to construct natural language instructions (Sanh et al., 2022; Wang et al., 2022f).
Apart from bimodal training and instruction finetuning, another recent trend in constructing code dataset is synthesizing data with powerful models such as ChatGPT. While this method is originally proposed for generating instruction data in natural language (Wang et al., 2023e; Honovich et al., 2023), Gunasekar et al. (2023) take one step further and synthesize 1B tokens of Python textbooks and coding exercises to pretrain a 1.3B model, achieving state-of-the-art results on HumanEval that’s comparable to much larger models trained on significantly larger datasets.
@@ 5.2 Encoders
Apart from these standard training objectives, many auxiliary objectives specifically designed for code have also been introduced. GraphCodeBERT (Guo et al., 2021a) and SynCoBERT (Wang et al., 2021c) both extract graphs from the source code (data flow graph and abstract syntax tree, respectively) and train the models to predict the typological relations between the nodes, while SynCoBERT and Code-MVP (Wang et al., 2022d) also add type inference to their pretraining stage in the form of tagging. Another common objective is contrastive learning: SynCoBERT and Code-MVP contrast between different views of the input (such as code, comment, AST, and transformed code), while DISCO (Ding et al., 2022a) constructs positive sample pairs by semantic-preserving transformations such as obfuscation, and negative pairs by injecting artificial bugs.
@@ 5.3 Encoder-Decoders
Inspired by these advantages of encoder-decoder architecture, many such models have been proposed for code processing. PyMT5 (Clement et al., 2020) and Mastropaolo et al. (2021) replicate the pretraining and multi-task finetuning process of T5 on code corpus, while Ahmad et al. (2021) introduce PLBART, a BART pretrained on 655GB combined data of Java, Pyhton, and natural language. Lachaux et al. (2021) argue that MLM could be too easy a task for programming languages as identifier names often occur multiple times in a single context window, and propose a deobfuscation pretraining objective, where the model is trained to convert obfuscated code back to its original form. Related to this method, we note that meaningful variable names have also been found to have a positive impact on the code generation process of large language models (Chen et al., 2023c).
Building on these early works, Wang et al. (2021e) propose CodeT5, which is pretrained alternatively with 1) T5’s original span corruption; 2) identifier tagging (where each token in the code input is tagged as either identifier or non-identifier); 3) masked identifier prediction (a special form of span corruption where all identifiers are masked); and 4) text-to-code & code-to-text generation. Its successor, CodeT5+ (Wang et al., 2023f), take inspiration from UL2 (Tay et al., 2023b) and introduce causal language modeling (CLM) into pretraining, along with additional contrastive objectives based on text-code matching.
AlphaCode (Li et al., 2022g) is also trained with multiple objectives, where the encoder is trained with MLM and the decoder is trained with CLM, with architecture modifications such as shallow-encoder & deep-decoder, multi-query attention (Shazeer, 2019), and being much larger than CodeT5 (up to 41B parameters). NatGen (Chakraborty et al., 2022a), on the other hand, is pretrained with a "naturalization" objective similar to deobfuscation: semantically equivalent but unnatural code is generated by predefined operations such as loop transformation, dead code injection, and variable renaming, and the model is pretrained to translate

### v6

@@ 2.1.4 Deployment and Operations
- Workflow Automation is the automation of workflows in Git. Due to the complexity involved in this task, it was only recently brought to attention after the advent of powerful LLMs.
2.1.5 LLM Evaluation
Apart from the previously covered stages in SE, we observe that programming-related evaluation of LLMs has been gaining momentum in both the NLP and the SE community since the release of Codex. Thus, we also list several novel tasks that appeared recently for this purpose. We discuss this topic in more detail in Section 7.
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=4.1em,font=,
where level=3text width=26em,font=,
where level=4text width=20.7em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[NLP in SE, fill=black, fill opacity=0.7, text=white
[Requirement, fill=yellow, fill opacity=0.5]
[Development, fill=Orange, fill opacity=0.2]
[Testing &
Analysis, fill=cyan, fill opacity=0.2]
[Deployment & Operations, fill=VioletRed, fill opacity=0.2
[LLM Evaluation, fill=Dandelion, fill opacity=0.4
[Code
Reasoning, fill=Dandelion, fill opacity=0.4
MMLU (Hendrycks et al., 2021b),
CodeQA (Liu & Wan, 2021),
CS1QA (Lee et al., 2022),
C-Eval (Huang et al., 2023b),
CMMLU (Li et al., 2023b),
CodeApex (Fu et al., 2023),
CRUXEval (Gu et al., 2024),
PythonIO (Zhang et al., 2024d)
, fill=Dandelion, fill opacity=0.4
[Coding for
Reasoning, fill=Dandelion, fill opacity=0.4
MathQA-Python (Austin et al., 2021),
Drori & Verma (2021),
Drori et al. (2022),
PAL (Gao et al., 2023b),
PoT (Chen et al., 2023e),
He-Yueya et al. (2023),
GSM8K-Python (Chowdhery et al., 2023; Wang et al., 2023h),
CoC (Li et al., 2023a),
Think-and-Execute (Chae et al., 2024)
, fill=Dandelion, fill opacity=0.4
[AIGC
Analysis, fill=Dandelion, fill opacity=0.4
[Vulnerabilities
, fill=Dandelion, fill opacity=0.4, text width=4em
Schuster et al. (2021),
Asare et al. (2023),
Sandoval et al. (2023),
Perry et al. (2023),
Hamer et al. (2024)
, fill=Dandelion, fill opacity=0.4
[Correctness
, fill=Dandelion, fill opacity=0.4, text width=4em
Nguyen & Nadi (2022),
Jesse et al. (2023),
Yetistiren et al. (2023),
Tambon et al. (2024)
, fill=Dandelion, fill opacity=0.4
[AIGC
Detection
, fill=Dandelion, fill opacity=0.4, text width=4em
DetectGPT4Code (Yang et al., 2023e),
CodeIP (Guan et al., 2024)
, fill=Dandelion, fill opacity=0.4
[Others
, fill=Dandelion, fill opacity=0.4, text width=4em
Understandability (Liu et al., 2023n),
Syntactic Robustness (Sarker et al., 2024),
Hallucinations (Liu et al., 2024),
Efficiency (Niu et al., 2024),
Data Leak (Wan et al., 2024)
, fill=Dandelion, fill opacity=0.4
- Code reasoning requires a model to reason about code or algorithms, and answer related questions which are written in multiple-choice format or free-form QA format, which may range from conceptual understanding to numerical calculation and complexity analysis. It often comes as a subset of composite evaluation benchmarks such as MMLU (Hendrycks et al., 2021b).
- Coding for reasoning is a special case of reasoning where LLMs solve complex reasoning problems by generating code that will be executed by external interpreters. This task abstracts the reasoning process from numerical calculations, and is thus of special interest in evaluating LLMs. More recently research has also shown that asking the LLM itself to simulate the execution of generated code can also help its reasoning process.
- AIGC analysis aims to analyze AI-generated code222In the AI community “AIGC” usually refers to AI-Generated Content. In this work we overload this expression with AI-Generated Code. from various aspects, such as correctness, vulnerabilities, and detection of AI-generated code.
@@ 2.1.7 Code LLMs for Low-Resource, Low-Level, and Domain-Specific Languages
This phenomenon also exists in code modeling, as most works listed in Figure 3-8 target the most popular programming languages such as C, Java, and Python. However, a major difference between NLP and SE is that in NLP, low-resource languages are often spoken by few people or even endangered, while in SE low-resource languages are often designed for specific domains and purposes, and thus have an active user community. Verilog, a hardware description language, is one such example, for which code modeling research is quite abundant (Thakur et al., 2023a; b; Lu et al., 2023b; Liu et al., 2023g; Tsai et al., 2023; Thorat et al., 2023; Liu et al., 2023h; Nadimi & Zheng, 2024).
From a different perspective, most of these popular languages are imperative languages, while few works have studied NLP and LLM applications in declarative or functional languages, except those concerning SQL and one recent work on Haskell (van Dam et al., 2024). However, research has shown that declarative languages are more aligned for optimization since they describe the “what” and not “how” in programming (Marcus et al., 2019; 2020), thus they are worth more attention in the future research.
@@ 2.2 Evaluation Metrics
As code generation models became more capable over the years, however, these metrics based on content-overlap have been found to be inadequate (Rozière et al., 2020; Hendrycks et al., 2021a; Austin et al., 2021), since functionally equivalent snippets of code can differ dramatically in their lexical forms. Consequently, researchers have turned their attention to functional correctness. One popular example of such metrics is pass@, proposed by Kulal et al. (2019) and refined by Chen et al. (2021b), which is an unbiased estimator of the model’s chance in passing all unit tests of a program with any of generated samples. This metric can be generalized to pass (Li et al., 2022h), which limits the number of model submissions to but allows filtering by unit tests given in the input from samples.
@@ 2.3 Program Synthesis
Many of the early methods for program synthesis are example-based (Menon et al., 2013), which means they induce programs from input-output examples, often in domain-specific languages (DSLs) such as FlashFill (Devlin et al., 2017a) and Karel333FlashFill is used in Microsoft Excel for string transformation. Karel is a simple programming language for educational purpose. (Bunel et al., 2018), as these languages are usually simple in syntax and structure.
As code generation models became more capable over the years, researchers started to pay attention to program synthesis in general-purpose programming languages as well. Hearthstone (Ling et al., 2016) and CONCODE (Iyer et al., 2018) are two of the early datasets, representing Python and Java respectively. Each example in Hearthstone is the description of a card in the game and its corresponding class implementation, while examples in CONCODE are simply Java methods paired with their natural-language documentation crawled from public GitHub repositories. Synthesizing programs from their corresponding natural language descriptions has since then become a standard practice in program synthesis, and has led to some of the most widely used benchmarks, such as HumanEval (Chen et al., 2021b), which has even been translated into multiple languages (Cassano et al., 2023b; Zheng et al., 2023a; Muennighoff et al., 2024). Some recent benchmarks use general-purpose languages but focus on specific domains, such as data science (Bavishi et al., 2019; Lai et al., 2023) or Jupyter notebooks (Agashe et al., 2019), while several math reasoning benchmarks have also been converted to programming tasks, including MathQA-Python (Amini et al., 2019; Austin et al., 2021) and GSM8K-Python (Cobbe et al., 2021; Chowdhery et al., 2023; Wang et al., 2023h).
@@ 2.4 Repository-Level Coding
Most tasks discussed in Section 2.1 are limited to a single file or even a single function, as cross-file code modeling poses challenges that are beyond the capability of most existing language models. Recently, however, position interpolation techniques (Chen et al., 2023d; Rozière et al., 2023; Peng et al., 2023a) have extended the context window of LLMs to hundreds of thousands of tokens, making it possible to contextualize coding activities within entire repositories. Several works have studied code completion (Shrivastava et al., 2023b; Ding et al., 2022b; Zhang et al., 2023b; Shrivastava et al., 2023a; Phan et al., 2024; Wu et al., 2024b) and generation (Liao et al., 2023; Zan et al., 2024) leveraging repository-level context, and corresponding benchmarks have been proposed Liu et al. (2023j); Ding et al. (2023); Li et al. (2024a). Bairi et al. (2023) investigate the more challenging tasks of repository-level API migration and temporal editing, while Jimenez et al. (2023) introduce a related benchmark, SWE-bench.
@@ 3.1 Causal Language Modeling
Unidirectional language models (also known as causal language models444The training objective of such language models is Causal Language Modeling (CLM), but also referred to as Next Token Prediction.) factor the probability of a sentence into the product of each token’s conditional probability with the chain rule. A piece of input text consisting of tokens is modeled as
@@ 3.3 Denoising Objectives
Tay et al. (2023b) argue that span corruption is also closely related to CLM, since one can mask out the whole input text as a single span and train the decoder to generate it autoregressively. Inspired by this relation, they propose UL2, which is the combination of many span corruption objectives that differ in corruption rate and span length. Applying it to both encoder-decoder models and decoder-only models, they find that encoder-decoder models perform better under the same computation budget constraint. Other researches have also found that such encoder-decoder models generally perform better than causal decoder-only models (Wang et al., 2022c; Soltan et al., 2022).
@@ 3.5 Implementation Design
As self-attention alone cannot capture the position information of each input token, the vanilla Transformer adds a non-learnable vector that’s dependent on the absolute position in the input sequence to the embedding of each token, which is known as cosine position embedding. Later works such as GPT and BERT use learnable embeddings instead, while T5 adopts relative position embedding (Shaw et al., 2018), where the information of relative position between query and key is injected in the computation of self-attention instead. A more recent scheme RoPE (Su et al., 2024) multiplies the keys and queries by a position-dependent rotation matrix, and is later shown to enable position interpolation for processing of longer sequences (Chen et al., 2023d; Rozière et al., 2023; Peng et al., 2023a). Alternatively, Press et al. (2022) propose ALiBi, which directly attenuates the attention scores according to the relative position between key and query. RoPE is popularized by its application in PaLM (Chowdhery et al., 2023) and GPT-NeoX (Black et al., 2022), while ALiBi is adopted by BLOOM (Scao et al., 2022). We refer to Zhao et al. (2023) for a survey on position embedding and interpolation in Transformers.
@@ 4 General Language Models for Code
Since language models scaled to hundreds of billions of parameters (Brown et al., 2020; Chowdhery et al., 2023), many of them have demonstrated non-trivial coding capability, even if they are not specifically designed or trained for code. Pioneered by Codex, researchers have also found continual pretraining on code to significantly benefit language models’ performance on code666While some works refer to this process as “finetuning on code”, it is still self-supervised in nature. Thus we choose to adopt the term “extra/additional/continual pretraining” in this work to avoid confusion with supervised in-task finetuning or instruction finetuning..
@@ 4.1 Off-the-Shelf Language Models
Another rising trend is mixture-of-expert models (Lepikhin et al., 2021; Fedus et al., 2022; Du et al., 2022). Notably, Jiang et al. (2024a) recently introduce Mixtral 8x7B, where only 13B parameters are activated for each token during inference, but achieving 40.2 on HumanEval and surpassing much larger dense models such as LLaMA 2. Similarly, Dai et al. (2024a) present DeepSeekMoE, where only 2.8B parameters are activated in a 16B model, scoring 26.8 on HumanEval.
@@ 5 Specialized Language Models for Code
As pretrained Transformers such as GPT and BERT achieved remarkable success in natural language processing, such model architectures, learning paradigms, and training objectives were soon adopted by the software engineering community to produce specialized models for code understanding and generation. In this section, we first review common datasets used for pretraining code language models (Section 5.1), and then dive into the complex family of code LMs by their model architecture: encoder-only models (Section 5.2), encoder-decoder models (Section 5.3), decoder-only models (Section 5.4), UniLM (Section 5.5),
@@
and diffusion models (Section 5.6). Lastly, in Section 5.7 we also illustrate the current trend of applying more recent techniques in NLP, such as instruction tuning (Wei et al., 2022b; Sanh et al., 2022; Chung et al., 2022) and reinforcement learning (Ouyang et al., 2022) to code processing. An overview of these pretrained models are provided in Table 3.
5.1 Training Dataset for Code
@@
Compared with natural language, another byproduct of scraping code from GitHub is commit histories, which consist of code before commit, code after commit, and a short message describing the commit, which can loosely serve as an instruction for language models. Muennighoff et al. (2024) utilize this feature and construct a 2GB dataset CommitPackFT containing 742K samples of instruction data for code, obviating the need of extensive human labor that’s required to construct natural language instructions (Sanh et al., 2022; Wang et al., 2022g).
Apart from bimodal training and instruction finetuning, another recent trend in constructing code dataset is synthesizing data with powerful models such as ChatGPT. While this method is originally proposed for generating instruction data in natural language (Wang et al., 2023g; Honovich et al., 2023), Gunasekar et al. (2023) take one step further and synthesize 1B tokens of Python textbooks and coding exercises to pretrain a 1.3B model, achieving state-of-the-art results on HumanEval that’s comparable to much larger models trained on significantly larger datasets.
@@ 5.2 Encoders
Apart from these standard training objectives, many auxiliary objectives specifically designed for code have also been introduced. GraphCodeBERT (Guo et al., 2021a) and SynCoBERT (Wang et al., 2021d) both extract graphs from the source code (data flow graph and abstract syntax tree, respectively) and train the models to predict the typological relations between the nodes, while SynCoBERT and Code-MVP (Wang et al., 2022e) also add type inference to their pretraining stage in the form of tagging. Another common objective is contrastive learning: SynCoBERT and Code-MVP contrast between different views of the input (such as code, comment, AST, and transformed code), while DISCO (Ding et al., 2022a) constructs positive sample pairs by semantic-preserving transformations such as obfuscation, and negative pairs by injecting artificial bugs.
@@ 5.3 Encoder-Decoders
Inspired by these advantages of encoder-decoder architecture, many such models have been proposed for code processing. PyMT5 (Clement et al., 2020) and Mastropaolo et al. (2021) replicate the pretraining and multi-task finetuning process of T5 on code corpus, while Ahmad et al. (2021) introduce PLBART, a BART pretrained on 655GB combined data of Java, Pyhton, and natural language. Lachaux et al. (2021) argue that MLM could be too easy a task for programming languages as identifier names often occur multiple times in a single context window, and propose a deobfuscation pretraining objective, where the model is trained to convert obfuscated code back to its original form. Related to this method, we note that meaningful variable names have also been found to have a positive impact on the code generation process of large language models (Chen et al., 2023e).
Building on these early works, Wang et al. (2021f) propose CodeT5, which is pretrained alternatively with 1) T5’s original span corruption; 2) identifier tagging (where each token in the code input is tagged as either identifier or non-identifier); 3) masked identifier prediction (a special form of span corruption where all identifiers are masked); and 4) text-to-code & code-to-text generation. Its successor, CodeT5+ (Wang et al., 2023h), take inspiration from UL2 (Tay et al., 2023b) and introduce causal language modeling (CLM) into pretraining, along with additional contrastive objectives based on text-code matching.
AlphaCode (Li et al., 2022h) is also trained with multiple objectives, where the encoder is trained with MLM and the decoder is trained with CLM, with architecture modifications such as shallow-encoder & deep-decoder, multi-query attention (Shazeer, 2019), and being much larger than CodeT5 (up to 41B parameters). NatGen (Chakraborty et al., 2022a), on the other hand, is pretrained with a "naturalization" objective similar to deobfuscation: semantically equivalent but unnatural code is generated by predefined operations such as loop transformation, dead code injection, and variable renaming, and the model is pretrained to translate